OpenStudy (anonymous):

roots for x^3-6x^2+7x-2 HELP

OpenStudy (anonymous):

please

OpenStudy (anonymous):

are they 1 and 2?

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

By trial, we can easily find that \(1\) is a root of the given polynomial, and therefore \(x-1\) is a factor of it. Using long division to divide \(x^3-6x^2+7x-2\) by \(x-1\), we find that \(x^3-6x^2+7x-2=(x-1)(x^2-5x+2)\). That means the polynomial has also the roots of the quadratic equation \(x^2-5x+2\), which are \(\frac{1}{2}(5+\sqrt{17})\), and \(\frac{1}{2}(5-\sqrt{17})\).

OpenStudy (anonymous):

gblover i answer this on your other q..lol